EpicLeo:

tan^2 x + sec^2 x / tan x sec x

Narad:

\[=\frac{\tan ^{2}x }{ tanx secx }+\frac{ \sec^{2} x}{ tanxsecx }\] \[=\frac{ tanx }{ secx }+\frac{ secx }{ tanx }\] \[=sinx+\frac{ 1 }{ sinx }\]